### v2

# Identifying faulty edges in resistive electrical networks

Abstract
Given a resistive electrical network, we would like to determine whether all the resistances (edges) in the network are working, and if not, identify which edge (or edges) is faulty. To make this determination, we are allowed to measure the effective resistance between certain pairs of nodes (which can be done by measuring the amount of current when one unit of voltage difference is applied at the chosen pair of nodes). The goal is to determine which edge, if any, is not working in the network using the smallest number of measurements. We prove rigorous upper and lower bounds on this optimal number of measurements for different classes of graphs. These bounds are tight for several of these classes showing that our measurement strategies are optimal.
Keywords: Electrical networks, Spectral graph theory, Information theory, Fault detection, Combinatorial optimization
1 Introduction
Diagnosing faults in electrical circuits and wiring is an important problem with applications in power systems, very large scale integration (VLSI), chip design, circuit board systems, to name a few. The literature is vast; the reader is referred to the survey [furse2020fault], a sample of more recent work [dwivedi2010fault, wang2019electromagnetic] and the references therein. There are a variety of approaches that are adopted in practice and new methods are regularly being invented even to the present. In this paper, we consider a particular formalization of this problem which, to the best of our knowledge, has not been studied rigorously in prior work. Nevertheless, we believe the model we study is natural, of practical relevance, and worth studying from a mathematical perspective. It is related to and inspired by work in detecting faults and structures in continuous conducting media; see the textbooks [holder2004electrical, bui2007fracture] and [adler2015electrical] for a more recent survey. Our problem can be seen as a discrete analogue of this problem from the literature on continuous media and related inverse problems. The resulting optimization problem is rich in its scope, spanning topics from combinatorial and spectral graph theory, integer and combinatorial optimization, information theory, and electrical network theory. Let us formally introduce the problem.
Definition 1.1.
[…]
Given two vertices $r,s\in V$, which are called the source and the sink terminals, the effective resistance $R_{rs}$ of the network between $r,s$ is defined as the reciprocal of the net current that would flow through the network if a unit voltage difference was applied to the source-sink pair.
The problem we wish to study is to detect whether all the resistance connections in the network are working as expected, or whether one of them has been damaged. To do this, we are allowed to probe the network by making current value measurements (or equivalently, effective resistance values) across pairs of terminals (vertices) and use these values to determine if any edge is faulty and if so, which one.
Definition 1.2.
[…]
A different version of the problem is obtained when the adversary changes the weight of $e^{*}$ to $\infty$ (that edge has been shorted, which is equivalent to contracting the edge).
One could allow for the possibility that no edge has been altered. Then, one has to detect if any edge has been altered and if so, identify which one. We give a short argument below that the number of measurements in this version of the problem is at most one more than the version in Definition 1.2 where a single, unknown edge has been altered. Consequently, we will focus on the version in Definition 1.2 for the remainder of the paper.
Let $\mathcal{A}\in\mathbb{R^{|T|\times|E|}}$ be the matrix where each row corresponds to a measurement $t\in T$ and each column corresponds to an edge $e\in E$ that has been altered. Each entry in $\mathcal{A}$ contains the effective resistance value for a given pair of source and sink terminals and a given altered edge. Consider the submatrix $\mathcal{A}_{S}\in\mathbb{R^{|S|\times|E|}}$ of values corresponding to measurements $S\subseteq T$ and all edges $E$. If all columns in $\mathcal{A}_{S}$ are different from one another, then the provided measurements $S$ satisfy the version of our problem given in Definition 1.2. On the other hand, if there are at least two columns that are the same, then the set $S$ of measurements will not distinguish between these two edges. Hence, an equivalent formulation of the problem is to find the smallest set $S\subseteq T$ of measurements such that $\mathcal{A}_{S}$ has distinct columns.
Suppose now we include a new column $c$ in $\overline{\mathcal{A}}=[\mathcal{A}|c]$ which contains the effective resistance values when no edge in $G$ has been altered. Consider any set $S$ of measurements such that $\mathcal{A}_{S}$ has distinct columns (i.e., $S$ is a set of measurements that can identify a faulty edge assuming one exists). If all columns of the updated submatrix $\overline{\mathcal{A}}_{S}$ are different from one another, then the measurements $S$ solve the problem including the possibility that no edge has been altered. Otherwise, there is one column in $\overline{\mathcal{A}}_{S}$ that is the same as column $c$ restricted to $\overline{\mathcal{A}}_{S}$ (there can be at most one such column since $\mathcal{A}_{S}$ has distinct columns). Let $e^{*}\in E$ be the edge corresponding to this column. If there is no measurement in $T$ such that the column corresponding to $e^{*}$ has a different entry compared to column $c$ in $\overline{\mathcal{A}}$, then the problem cannot be solved using measurements in $T$. In other words, no measurement is able to distinguish between the possibility that $e^{*}$ has been altered and the possibility that no edge has been altered. Thus, we may assume there exists a measurement $s^{*}\in T$ such that the corresponding entry of column of $c$ will be different from that associated with $e^{*}$. As a result, $S^{*}=S\cup\{s^{*}\}$ is a set of measurements that will solve the problem with the additional possibility that no edge has been altered. Since $|S^{*}|=|S|+1$, at most one additional measurement is required if the possibility of no edge being altered is included in the problem.
We also note that in Definition 1.2, the measurements $S\subseteq T$ are made non-adaptively, i.e., the result of a measurement is not used to decide what the next measurement is. One must decide on the set $S$ of measurements up front and then the observed measurement values must be used to identify the faulty edge. Our results below apply to this non-adaptive setting. The analysis of adaptive measurement strategies is left for future work.
1.1 Our contributions
We provide tight upper and lower bounds on the smallest number of measurements for the fault detection problem in various families of electrical networks. In particular, we are able to fully resolve the problem for complete graphs and complete $k$-partite graphs for $k\geq 2$. The formal statements are in Section 2 and their proofs are provided in Section LABEL:sec:proofs. The proof techniques draw upon combinatorial and spectral graph theory, electrical network theory, and ideas from information theory. We note briefly that each effective resistance measurement provides a real number; thus, a single measurement potentially provides infinitely many bits of information. For this reason, classical information theoretic lower bound arguments based on bits of information do not immediately apply. While our final results are mathematically precise and rigorous, our search for the best measurement strategy in each of these cases involved modeling the problem as a set covering/integer programming problem and using computational experiments to guide us towards the optimal strategy for measurements. Thus, the problem involves a nice interplay between rigorous mathematical analysis and computer-aided search using discrete optimization techniques. More broadly, we believe that the fault detection problem we set up in Definition 1.2 is a new, challenging combinatorial optimization problem that is rich in its mathematical structure and is motivated by a fundamental problem in fault diagnosis in electrical networks, which has diverse applications. We close the paper in Section LABEL:sec:future_avenues with some future directions, including some concrete open questions, that we hope can spawn new and interesting research avenues for the discrete mathematics and optimization community.
1.2 Related literature
As mentioned before, the formalization of fault detection given in Definition 1.2 does not seem to have received a great deal of attention in terms of obtaining provable bounds on the smallest number of measurements. Nevertheless, there are several very related strands of work on fault detection in electrical networks that we review next, which place our work in the broader context of fault diagnosis in electrical systems.
The closest line of work is the investigation in [kahng1998test], which is motivated by a fault detection problem in VLSI systems. In this paper, the authors consider the problem of testing whether all edges in a tree network are working properly; thus, the problem is called the tree testing problem. In the language of Definition 1.2, the graph $G$ is a tree, all edge weights are $1$, and the goal is to simply detect if any edge has been altered or not, without necessarily pinpointing which one if a fault is indeed detected.
A very closely related line of work originates in the literature on inverse problems, especially in the so-called Calderón problem. The discrete version of the Calderón problem is to recover the values of all resistors in a resistive network given a possible set of measurements that one can make. While there is no adversary altering the resistances as in Definition 1.2, obviously if one can determine all the resistance values, then one can simply check if any one of them differs from the original value. Thus, the discrete Calderón problem requires one to deduce much more information from the measurements, compared to the faulty edge detection problem in Definition 1.2 (and therefore may require a much larger number of measurements). A good introduction to the discrete Calderón problem is the book [curtis2000inverse], which also surveys relevant prior work.  [kazakov2025inverse] is a recent paper that surveys work in the past couple of decades and provides some new mathematical foundations for the problem. The original Calderón problem is defined for a continuous conducting medium as opposed to a discrete electrical network modeled by a graph. One is allowed to make current and voltage measurements on the boundary of a connected domain $\Omega\subseteq\mathbb{R}^{d}$ and from this one has to infer the electrical properties in the interior of $\Omega$. The problem forms the mathematical foundation for Electrical Impedance Tomography (EIT) [holder2004electrical, adler2015electrical, borcea2002electrical], which has numerous applications ranging from computational medicine, geophysics, to testing for structural defects in solids. Much of the literature uses the terms EIT and Calderón problem interchangeably. The tools and techniques are quite different from the discrete setting (and especially the techniques we use in this paper), drawing upon methods from PDEs and classical inverse problems. A continuous version of the problem we study in this paper (Definition 1.2) has been studied, arising out of the work on the continuous Calderón problem. Here, one considers a fracture or a crack in the conducting medium, modeled by a curve that is perfectly insulating. The goal is to, again, use boundary measurements to determine the location/shape of the curve. The problem was introduced in a seminal paper by Friedman and Vogelius [friedman1989determining], with several follow up works; see [alessandrini1993stable, alessandrini1995stability, alessandrini1996unique, alessandrini1997determining, santosa1991computational] for a sample and the textbook [bui2007fracture] for more details.
In [shi1999diagnosis, shi2001structural], the authors consider the problem of detecting faults in wiring networks. They build on prior work which can be traced back to a seminal paper by Kautz [kautz1974testing]. Here, the vertices of a graph model different logical units (often called nets) in a circuit board and each edge denotes a possible site of a “fault”. One is allowed to probe the logical units and use the responses to detect these faults, if any. The mathematical formalization is the following: given a (simple) graph $G=(V,E)$, one can query a subset $U\subseteq V$ and the response is a subset $Q(U)$ which is the set of all vertices that can be reached by a path from a vertex in $U$. The goal is to find the smallest set of queries to identify all the connected components of $G$ (the vertices in the connected components of size two or more correspond to the logical units that are connected by faulty edges). A group testing [aldridge2019group] approach to this problem was considered in [chen1989detecting] and is discussed in the papers [shi1999diagnosis, shi2001structural] cited above, along with other relevant work on this problem.
2 Statement of results
We consider the fault detection problem in Definition 1.2 where all the unaltered weights are $1$, i.e., the network is made of unit resistance (conductance) edges; we use $\mathbf{1}$ to denote the vector of all ones. Moreover, we allow all possible measurements, i.e., $T=V\times V$ (see Definition 1.2). We present upper and lower bounds when the electrical network’s underlying graph is a complete graph or a complete $k$-partite graph.
Complete graphs
The following theorem gives the number of measurements
needed for complete graphs (i.e., graphs that have an edge between every pair of vertices).
Theorem 2.1.
The smallest number of measurements needed to solve the faulty edge detection problem in a complete graph on $n\geq 6$ vertices is exactly $\left\lceil\frac{2n}{3}\right\rceil$.
Complete $k$-partite graphs
A complete $k$-partite graph is one where the set of vertices can be partitioned into $k$ independent sets and there is an edge between every pair of vertices in different partitions. We first state exact bounds for the bipartite ($k=2$) and tripartite ($k=3$) cases, and then state a generalization for arbitrary $k\geq 2$.
Theorem 2.2.
Let $(V,E,\mathbf{1})$ be an electrical network such that the underlying graph $G=(V,E)$ is a complete bipartite graph with $|V|=n$ and partitions $p_{\beta},p_{\gamma}$ with $|p_{\beta}|\leq|p_{\gamma}|$.
When $|p_{\beta}|<|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor{\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|}\right\rfloor-1$ when $|p_{\gamma}|-|p_{\beta}|\equiv 0\mod{3}$, and equal to $\left\lfloor\frac{2}{3}|p_{\gamma}|+\frac{1}{3}|p_{\beta}|\right\rfloor$ when $|p_{\gamma}|-|p_{\beta}|\equiv(1\textrm{ or }2)\mod{3}$.
When $|p_{\beta}|=|p_{\gamma}|$, the smallest number of measurements needed to solve the faulty edge detection problem is $\left\lfloor\frac{2n}{3}\right\rfloor-1$ when $|p_{\beta}|\equiv 0\textrm{ or }1\mod{3}$, and equal to $\left\lfloor\frac{2n}{3}\right\rfloor$ when $|p_{\beta}|\equiv 2\mod{3}$.
Theorem 2.3.
Let $(V,E,\mathbf{1})$ be an electrical network such that the underlying graph $G=(V,E)$ is a tripartite graph with $|V|=n$ and partitions $p_{\alpha},p_{\beta},p_{\gamma}$ with $|p_{\alpha}|\leq|p_{\beta}|\leq|p_{\gamma}|$. The smallest number of measurements needed to solve the faulty edge detection problem is given in Table LABEL:tab:bounds_tripart.
Partition Sizes Upper Bound Lower Bound
$|p_{\alpha}|<|p_{\beta}|<|p_{\gamma}|$ $\left\lceil{\frac{n-3}{2}}\right\rceil$ $\left\lceil\frac{n-3}{2}\right\rceil$
$|p_{\alpha}|=|p_{\beta}|<|p_{\gamma}|$